OpenStudy (anonymous):

find the domain and range the verticle and horizaontal asymptotes of the graph.... f(x)=x^2-2x-3

OpenStudy (anonymous):

@uri

OpenStudy (sidsiddhartha):

ok first domain: \[f(x)=x^2-2x-3\] its a simple quadratic equation so domain is R(all real numbres)

OpenStudy (sidsiddhartha):

now for range we write\[f(x)=x^2-2x-3=(x^2-2.x.1+1^2)-4=(x-1)^2-4\] so range:\[f(x) \in \mathbb{R} :f(x) \ge -4\]

OpenStudy (sidsiddhartha):

and as its a quadratic equation so theres no asymptotes okay ??